Please can someone help. Thank You.
Solve the equation. Be sure to show all of your work.
-14a + 3 = 10

Answer by Alan3354(69443)

Solve the equation. Be sure to show all of your work.
-14a + 3 = 10
----------------
Subtract 3
-14a = 7
Divide by 14
-a = 1/2
Divide by -1
a = -1/2